### v2

# Tracking Brownian fluid particles in large eddy simulations

Abstract
In this paper, we establish a numerical method for simulation of wall-bounded incompressible turbulent flows by integrating the technology of random vortex method with the core idea of Large Eddy Simulation (LES). Specifically, we utilize the filtering function in LES, interpreted as spatial averaging, along with the integral representation theorem for parabolic equations, to achieve a closure numerical scheme which may be used for calculating solutions of Navier-Stokes equations. This approach circumvents the challenge associated with handling the non-locally integrable 3-dimensional integral kernel in the random vortex method and facilitates the computation of numerical solutions for flow systems via Monte-Carlo method. Comprehensive numerical simulations are carried out for turbulent and laminar flows in full space and wall-bounded space, considering both two-dimensional and three-dimensional cases, thereby demonstrating the validity and effectiveness of the method.
Key words: large eddy simulation, random vortex method, incompressible fluid flow, turbulence simulation
MSC classifications: 76M35, 76M23, 60H30, 65C05, 68Q10,
1 Introduction
The main goal of the present paper is to propose a numerical scheme for simulating wall-bounded viscous and incompressible fluid flows, based on a stochastic integral representation of solutions to linear parabolic equations (see Section 5 below) to be established in the present paper, together with a key idea from the large eddy simulation (LES, cf. [16], [10] and [5], also cf. the monographs [1] and [9]). The method also applies to ideal models for viscous incompressible fluid flows in the whole space. LES is a numerical method extensively used for simulating turbulent flows (cf. [9] and [20]). In this article, we introduce a natural closure to LES models by using a stochastic formulation of viscous incompressible fluid flows in terms of Brownian fluid particles.
An incompressible viscous flow past a flat plate is modeled with a time dependent velocity
$u(x,t)=(u^{1}(x,t),\cdots,u^{d}(x,t))$
in $D:=\mathbb{R}_{+}^{d}=\{x:x\in\mathbb{R}^{d},x_{d}>0\}$ (where $d=2$ or $3$), so its boundary where $x_{d}=0$ models a solid wall. The fluid density $\rho(x,t)$ and pressure $p(x,t)$ are other fluid flow quantities in this model. Due to the incompressibility of the fluid flow, $\rho(x,t)$ stays as a constant, thus, without losing generality, we may assume that $\rho(x,t)=1$ for all $x$ and $t$. The dynamics of the fluid flow is described by the Navier-Stokes equations
$\frac{\partial}{\partial t}u^{i}+(u\cdot\nabla)u^{i}=\nu\Delta u^{i}-\frac{\partial}{\partial x_{i}}p+F^{i}$ (1)
for $i=1,\ldots,d$, and
$\nabla\cdot u=0$ (2)
in $D\times[0,\infty)$, subject to the non-slip condition that $u(x,t)=0$ for $x\in\partial D$ and $t>0$. Here $\nu>0$ is the kinematic viscosity, and $F(x,t)=(F^{1}(x,t),\cdots,F^{d}(x,t))$ stands for an external force applying on the fluid flow. With increasing computational power, numerical methods like finite difference, finite element method, fast Fourier transform and etc. may be applied to calculating numerical solutions of (1, 2).
If a viscous fluid flow moves along a solid wall with a high speed, then substantial vorticity may be created at the boundary which in turn generates turbulent motion in a thin layer next to the boundary, and may lead to flow separation from the wall. The small-scale energy dissipation in turbulent motion may be described by Kolmogorov’s theory, known as the K41 theory, while the global motion of boundary turbulence may be calculated by means of LES, or other numerical schemes.
The present work aims to propose a method for implementing LES by taking advantage of a stochastic integral representation of the velocity vector field of incompressible viscous fluid flows past a solid wall.
To motivate our approach, let us review the main ideas in the well-known vortex method (cf. [3], [4] and [11]). In the random vortex method, the main fluid dynamical variable is the vorticity, $\omega=\nabla\wedge u$, whose dynamics is described by the vorticity transport equation
$\frac{\partial}{\partial t}\omega^{i}+(u\cdot\nabla)\omega^{i}=\nu\Delta\omega^{i}+(\omega\cdot\nabla)u^{i}+(\nabla\wedge F)^{i}$ (3)
where $\omega^{i}=\varepsilon^{ijk}\frac{\partial}{\partial x_{j}}u^{k}$ for $i=1,2,3$. The flow velocity $u$ is recovered by solving the Poisson equation that $\Delta u=-\nabla\wedge\omega$.
It is appropriate to set up a convention about two-dimensional (2D) fluid flows. By a 2D flow we mean a 3D flow with velocity $u=(u^{1},u^{2},u^{3})$ where $u^{3}=0$, and $u^{1}$, $u^{2}$ depend only on coordinates $x_{1}$, $x_{2}$. Hence, for a 2D flow, $\omega^{1}=0$, $\omega^{2}=0$ identically, $\omega^{3}=\frac{\partial}{\partial x_{1}}u^{2}-\frac{\partial}{\partial x_{2}}u^{1}$, and $\omega$ is identified as a scalar function – its third component $\omega^{3}$, unless otherwise indicated. In particular the vorticity stretching term $(\omega\cdot\nabla)u$ vanishes identically for a 2D flow, and the vorticity equation becomes
$\frac{\partial}{\partial t}\omega+(u\cdot\nabla)\omega=\nu\Delta\omega+\nabla\wedge F$
where $\nabla\wedge F=\frac{\partial}{\partial x_{1}}F^{2}-\frac{\partial}{\partial x_{2}}F^{1}$ in this case. Hence, the vorticity $\omega$ can be represented in terms of the fundamental solution $p(s,\xi,t,y)$ of the parabolic operator $L_{-u}-\frac{\partial}{\partial t}$ (where we have introduced a notation that $L_{-u}=\nu\Delta-u\cdot\nabla$), and the initial vorticity $\omega_{0}$. Indeed
$\omega(x,t)=\int_{\mathbb{R}^{2}}p(0,\xi,t,y)\omega_{0}(\xi)\textrm{d}\xi+\int_{0}^{t}\int_{\mathbb{R}^{2}}p(s,\xi,t,y)\nabla\wedge F(\xi,s)\textrm{d}\xi\textrm{d}s.$ (4)
Since $\nabla\cdot u=0$, $p(s,\xi,t,y)$ is the transition probability density of the Brownian fluid particles $X$ with velocity $u$. By combining with the Biot-Savart law, the previous integral representation yields an integral representation for the velocity
$u(x,t)=\int_{\mathbb{R}^{2}}\mathbb{E}\left[K(x,X_{t}^{0,\xi})\omega_{0}(\xi)\right]\textrm{d}\xi+\int_{0}^{t}\int_{\mathbb{R}^{2}}\mathbb{E}\left[K(x,X_{t}^{s,\xi})\nabla\wedge F(\xi,s)\right]\textrm{d}\xi\textrm{d}s,$ (5)
where $X^{s,\xi}$ denotes the Taylor diffusion initiated from location $\xi$ at instance $s\geq 0$, that is, the trajectories of Brownian fluid particles with velocity $u$ determined by Itô’s stochastic differential equations
$\textrm{d}X_{t}^{s,\xi}=u(X_{t}^{s,\xi},t)\textrm{d}t+\sqrt{2\nu}\textrm{d}B_{t},\quad X_{s}^{s,\xi}=\xi,$ (6)
where $B_{t}(t\geq s)$ is Brownian motion on a probability space. $X$ is a diffusion process with infinitesimal generator $L_{u}$ (which is the $L^{2}$-adjoint operator of $L_{-u}$ as $\nabla\cdot u=0$).
Two equations (5, 6) consist of a closed McKean-Valsov type stochastic differential equations, which in turn may be used in the design of numerical schemes for calculating the flow velocity. It is possible to formulate the random vortex scheme for three-dimensional viscous flows, which, however, involves an integral kernel which is not locally integrable, leading to difficulties in numerical computation of 3D flows. See more details and related work in [7] and [14].
It is also possible to derive a random vortex formulation for flows constrained on a finite region and for flows past a solid wall. Although the random vortex method is mathematically elegant, it is effective only for certain types of flows and is often computationally expensive. For simulations of turbulent flows, schemes for calculating solutions of turbulent flows from their dynamical equations at separate space scales shall be advocated. At small scale, the universality principle of turbulence implies that the statistical properties of turbulence, such as energy dissipation and the structure of velocity gradients, become independent of the large-scale features of the flow and are governed by universal scaling laws. This motivates us to develop a new random vortex method that focuses more on relatively large scale flows.
The main contribution of the present work is that the integral representation of the fluid flow in terms of Brownian fluid particles gives rise to a natural closure of the averaged Navier-Stokes equations without further modelling the stress tensor $\tau$.
The remainder of the paper is organized as follows. In Section 2, the theoretical foundation of the random LES, proposed in this work, will be established including the details of the formulation of incompressible viscous fluid flows in terms of Brownian fluid particles. In Section 3, the random LES flows past a wall is established. The random LES are formulated in terms of stochastic differential equations involving the distribution of Brownian fluid particles. Various numerical experiments based on random LES are reported in Section 4. Finally, in an appendix, Section 5, we include the necessary results on the functional integral representation for solutions of linear parabolic systems, used in the formulation of the random LES.
2 Theoretical aspect of random LES
Let us first reformulate the fluid dynamics in terms of Brownian fluid particles, a device firstly put forward by G. Taylor [19]. Although it has not been established rigorously yet, we will nevertheless assume that the velocity $u(x,t)$ is bounded and smooth in $D\times[0,T)$ (for some $T>0$) and at least $C^{1}$ up to the boundary $\partial D$. As a consequence, $p(x,t)$ shares the same regularity as that of $u(x,t)$. The dynamics of the flow is determined by the Navier-Stokes equations (1, 2) together with the non-slip boundary condition. We introduce the theoretical framework of random LES.
2.1 Computation of the pressure gradient
Observe that there are two non-linear terms appearing in the Navier-Stokes equations (1), the non-linear convection term $(u\cdot\nabla)u$ and the pressure gradient $\nabla p$. In all numerical methods, one has to deal with the pressure gradient $\nabla p$. In Direct Numerical Simulation (DNS), one can select an appropriate basis, and project the Navier-Stokes equations to the space of divergence-free vector fields for eliminating the pressure gradient $\nabla p$. However, this method works well only for special class of regions. In random vortex method, one instead takes the vorticity $\omega=\nabla\wedge u$ to be the main fluid dynamical variable, and projects the Navier-Stokes equation to the differential of $u$, which is the vorticity transport equation. One can represent the vorticity $\omega$ in terms of Brownian fluid particles $X$. This also allows for the application of finite difference or finite element methods to both the vorticity transport equation and the Poisson equation.
The method we are going to propose, which combines the idea of Brownian fluid particles with an idea from LES, is based on a functional integral representation of solutions to linear parabolic equations in terms of distributions of Brownian fluid particles, cf. Section 5. Unlike the random vortex approach, although still utilizing the distribution of the Brownian fluid particles, we calculate the velocity directly instead of dealing with the vorticity equation. Hence we need to resolve the pressure gradient $\nabla p$ first.
Let $D=\mathbb{R}_{+}^{d}$ (where $d=2$ or $3$) whose boundary $\partial D$ is described by the equation that $x_{d}=0$, and is identified with $\mathbb{R}^{d-1}$. The unit normal pointing outwards to $\partial D$ is a constant vector $\boldsymbol{n}=(0,\cdots,0,-1)$, and the normal derivative $\frac{\partial}{\partial\boldsymbol{n}}=\left.-\frac{\partial}{\partial x_{d}}\right|_{x_{d}=0}$.
Since $\nabla\cdot u=0$, the pressure $p(x,t)$ at every instance $t>0$ is recovered by solving the Poisson equation
$\Delta p=-\frac{\partial u^{j}}{\partial x_{i}}\frac{\partial u^{i}}{\partial x_{j}}+\nabla\cdot F.$ (7)
Suppose the Navier-Stokes equations continue to be valid up to the boundary, then the pressure $p$ at the boundary $\partial D$ is subject to the Neumann boundary condition that
$\frac{\partial}{\partial\boldsymbol{n}}p=\nu\left(\frac{\partial}{\partial x_{1}}\frac{\partial u^{1}}{\partial x_{3}}+\frac{\partial}{\partial x_{2}}\frac{\partial u^{2}}{\partial x_{3}}\right)-F^{3}\quad\textrm{ on }\partial D$ (8)
where the right-hand side is evaluated along $\partial D$. Since $\frac{\partial}{\partial\boldsymbol{n}}=\left.-\frac{\partial}{\partial x_{3}}\right|_{x_{3}=0}$, the boundary condition may be written as
$\left.\frac{\partial p}{\partial x_{3}}\right|_{x_{3}=0}=-\nu\left(\frac{\partial}{\partial x_{1}}\frac{\partial u^{1}}{\partial x_{3}}+\frac{\partial}{\partial x_{2}}\frac{\partial u^{2}}{\partial x_{3}}\right)+F^{3}\quad\textrm{ on }x_{3}=0.$ (9)
Similarly for 2D flows
$\left.\frac{\partial p}{\partial x_{2}}\right|_{x_{2}=0}=-\nu\frac{\partial}{\partial x_{1}}\frac{\partial u^{1}}{\partial x_{2}}+F^{2}\quad\textrm{ on }x_{2}=0.$ (10)
We shall now recall the integral representation for solutions to the boundary problem (9) of the Poisson equation (7). According to the reflection principle, the Green function for $\mathbb{R}_{+}^{d}$, subject to the Neumann boundary condition, denoted by $H_{+}^{d}(x,y)$, is given by the following formula
$H_{d}^{+}(x,y)=\Gamma_{d}(x-y)+\Gamma_{d}(x-\bar{y})\quad\textrm{ for }x\neq y,\bar{y},$ (11)
where $\Gamma_{d}$ is the elementary solution to the Laplace equation on $\mathbb{R}^{d}$:
$\Gamma_{2}(x)=\frac{1}{2\pi}\ln|x|\quad\textrm{ and }\Gamma_{d}(x)=-\frac{1}{(d-2)s_{d-1}}\frac{1}{|x|^{d-2}}\textrm{ for }x\neq 0,$ (12)
for $d\geq 3$, where $s_{d-1}$ is the area of $d-1$ dimensional sphere, $s_{1}=2\pi$ and $s_{2}=4\pi$.
The Green formula yields the following representation (for simplicity the time variable $t$ in $p(x,t)$ is suppressed)
$p(x)$ $=\int_{\mathbb{R}_{+}^{d}}H_{d}^{+}(x,y)\Delta p(y)\textrm{d}y-\int_{\left\{y_{d}=0\right\}}H_{d}^{+}(x,y)\frac{\partial p}{\partial\boldsymbol{n}}(y)\textrm{d}y_{1}\cdots\textrm{d}y_{d-1}$
$=\int_{\mathbb{R}_{+}^{d}}H_{d}^{+}(x,y)\Delta p(y)\textrm{d}y$
$\;+\int_{\mathbb{R}^{d-1}}H_{d}^{+}(x,(y_{1},\cdots,y_{d-1},0))\left.\frac{\partial p}{\partial y_{d}}\right|_{y_{d}=0}(y_{1},\cdots,y_{d-1},0)\textrm{d}y_{1}\cdots\textrm{d}y_{d-1}$ (13)
for $x\in\mathbb{R}_{+}^{d}$. Noticing that we are only interested in the gradient of pressure $\nabla p$, which can be calculated therefore by the following formula:
$\nabla p(x)$ $=\int_{\mathbb{R}_{+}^{d}}K_{d}^{+}(x,y)\Delta p(y)\textrm{d}y$
$\;+\int_{\mathbb{R}^{d-1}}K_{d}^{+}(x,(y_{1},\cdots,y_{d-1},0))\left.\frac{\partial p}{\partial y_{d}}\right|_{y_{d}=0}(y_{1},\cdots,y_{d-1},0)\textrm{d}y_{1}\cdots\textrm{d}y_{d-1},$ (14)
where $K_{d}^{+}(x,y)=\nabla_{x}H_{d}^{+}(x,y)$, the Biot-Savart kernel for the half space $\mathbb{R}_{+}^{d}$.
If $d=2$, then
$K_{2}^{+}(x,y)=\frac{1}{2\pi}\left(\frac{x-y}{|x-y|^{2}}+\frac{x-\bar{y}}{|x-\bar{y}|^{2}}\right)\quad\textrm{ for }x\neq y,\bar{y},$ (15)
and
$K_{2}^{+}((x_{1},x_{2}),(y_{1},0))=\frac{1}{\pi}\frac{(x_{1}-y_{1},x_{2})}{|x_{1}-y_{1}|^{2}+|x_{2}|^{2}}\quad\textrm{ for }x_{2}\neq 0.$ (16)
If $d=3$, then
$K_{3}^{+}(x,y)=\frac{1}{4\pi}\left(\frac{x-y}{|x-y|^{3}}+\frac{x-\bar{y}}{|x-\bar{y}|^{3}}\right)\quad\textrm{ for }x\neq y,\bar{y},$ (17)
and
$K_{3}^{+}((x_{1},x_{2},x_{3}),(y_{1},y_{2},0))=\frac{1}{2\pi}\frac{\left(x_{1}-y_{1},x_{2}-y_{2},x_{3}\right)}{\left(|x_{1}-y_{1}|^{2}+|x_{2}-y_{2}|^{2}+|x_{3}|^{2}\right)^{\frac{3}{2}}}$ (18)
for $x_{3}\neq 0$.
Lemma 1.
(2D flows). For a 2D incompressible viscous flow the gradient of the pressure at any instance has the following singular integral representation
$\nabla p(x)$ $=\int_{\mathbb{R}_{+}^{2}}K_{2}^{+}(x,y)\Delta p(y)\textrm{d}y+\int_{\mathbb{R}}K_{2}^{+}(x,(y_{1},0))F^{2}(y_{1},0)\textrm{d}y_{1}$
$\;+\nu\int_{\mathbb{R}}\frac{\partial K_{2}^{+}}{\partial y_{1}}(x,(y_{1},0))\frac{\partial u^{1}}{\partial y_{2}}(y_{1},0)\textrm{d}y_{1}$ (19)
for $x\in\mathbb{R}_{+}^{2}$, where the last singular integral kernel is given by
$\frac{\partial K_{2}^{+}}{\partial y_{1}}((x_{1},x_{2}),(y_{1},0))=\frac{1}{\pi}\left(\frac{|x_{1}-y_{1}|^{2}-|x_{2}|^{2}}{\left(|x_{1}-y_{1}|^{2}+|x_{2}|^{2}\right)^{2}},\frac{2x_{2}(x_{1}-y_{1})}{\left(|x_{1}-y_{1}|^{2}+|x_{2}|^{2}\right)^{2}}\right)$ (20)
for $x_{2}>0$.
Proof.
By using the integral representation
$\nabla p(x)$ $=\int_{\mathbb{R}_{+}^{2}}K_{2}^{+}(x,y)\Delta p(y)\textrm{d}y$
$\;+\int_{\mathbb{R}^{1}}K_{2}^{+}(x,(y_{1},0))\left(-\nu\frac{\partial}{\partial y_{1}}\frac{\partial u^{1}}{\partial y_{2}}+F^{2}\right)(y_{1},0)\textrm{d}y_{1}.$
For the integral involving the iterated derivative of $u$, we may perform integration by parts we obtain that
$\nabla p(x)$ $=\int_{\mathbb{R}_{+}^{2}}K_{2}^{+}(x,y)\Delta p(y)\textrm{d}y+\int_{\mathbb{R}^{1}}K_{2}^{+}(x,(y_{1},0))F^{2}(y_{1},0)\textrm{d}y_{1}$
$\;+\nu\int_{\mathbb{R}^{1}}\frac{\partial K_{2}^{+}}{\partial y_{1}}(x,(y_{1},0))\frac{\partial u^{1}}{\partial y_{2}}(y_{1},0)\textrm{d}y_{1}$
where the kernel for the third singular integral can be worked out as
$\frac{\partial K_{2}^{+}}{\partial y_{1}}((x_{1},x_{2}),(y_{1},0))$ $=\frac{1}{\pi}\frac{(-1,0)}{|x_{1}-y_{1}|^{2}+|x_{2}|^{2}}+\frac{2}{\pi}\frac{\left(|x_{1}-y_{1}|^{2},x_{2}(x_{1}-y_{1})\right)}{\left(|x_{1}-y_{1}|^{2}+|x_{2}|^{2}\right)^{2}}$
$=\frac{1}{\pi}\left(\frac{|x_{1}-y_{1}|^{2}-|x_{2}|^{2}}{\left(|x_{1}-y_{1}|^{2}+|x_{2}|^{2}\right)^{2}},\frac{2x_{2}(x_{1}-y_{1})}{\left(|x_{1}-y_{1}|^{2}+|x_{2}|^{2}\right)^{2}}\right).$
By utilising the same argument, we may obtain a similar representation theorem stated as the following.
Lemma 2.
(3D flows) For 3D flows in $\mathbb{R}_{+}^{3}$ the gradient of the pressure is represented as
$\nabla p(x)$ $=\int_{\mathbb{R}_{+}^{3}}K_{3}^{+}(x,y)\Delta p(y)\textrm{d}y+\int_{\mathbb{R}^{2}}K_{3}^{+}(x,(y_{1},y_{2},0))F^{3}(y_{1},y_{2},0)\textrm{d}y_{1}\textrm{d}y_{2}$
$+\nu\int_{\mathbb{R}^{2}}\frac{\partial K_{3}^{+}}{\partial y_{1}}(x,(y_{1},y_{2},0))\frac{\partial u^{1}}{\partial y_{3}}(y_{1},y_{2},0)\textrm{d}y_{1}\textrm{d}y_{2}$
$+\nu\int_{\mathbb{R}^{2}}\frac{\partial K_{3}^{+}}{\partial y_{2}}(x,(y_{1},y_{2},0))\frac{\partial u^{2}}{\partial y_{3}}(y_{1},y_{2},0)\textrm{d}y_{1}\textrm{d}y_{2},$ (21)
where
$\frac{\partial K_{3}^{+}}{\partial y_{1}}(x,y)=\frac{\left(2|x_{1}-y_{1}|^{2}-|x_{2}-y_{2}|^{2}-|x_{3}|^{2},3(x_{2}-y_{2})(x_{1}-y_{1}),3x_{3}(x_{1}-y_{1})\right)}{2\pi\left(|x_{1}-y_{1}|^{2}+|x_{2}-y_{2}|^{2}+|x_{3}|^{2}\right)^{\frac{5}{2}}}$ (22)
and
$\frac{\partial K_{3}^{+}}{\partial y_{2}}(x,y)=\frac{\left(3(x_{1}-y_{1})(x_{2}-y_{2}),2|x_{2}-y_{2}|^{2}-|x_{1}-y_{1}|^{2}-|x_{3}|^{2},3x_{3}(x_{2}-y_{2})\right)}{2\pi\left(|x_{1}-y_{1}|^{2}+|x_{2}-y_{2}|^{2}+|x_{3}|^{2}\right)^{\frac{5}{2}}}$ (23)
for $x=(x_{1},x_{2},x_{3})$ with $x_{3}>0$ and $y=(y_{1},y_{2},0)$.
Proof.
Indeed, by the Green formula, and using the boundary condition for $p$, we obtain
$\nabla p(x)$ $=\int_{\mathbb{R}_{+}^{3}}K_{3}^{+}(x,y)\Delta p(y)\textrm{d}y+\int_{\mathbb{R}^{2}}K_{3}^{+}(x,(y_{1},y_{2},0))\left.\frac{\partial p}{\partial y_{3}}\right|_{y_{3}=0}(y_{1},y_{2},0)\textrm{d}y_{1}\textrm{d}y_{2}$
$\;=\int_{\mathbb{R}_{+}^{3}}K_{3}^{+}(x,y)\Delta p(y)\textrm{d}y+\int_{\mathbb{R}^{2}}K_{3}^{+}(x,(y_{1},y_{2},0))F^{3}(y_{1},y_{2},0)\textrm{d}y_{1}\textrm{d}y_{2}$
$\;+\int_{\mathbb{R}^{2}}K_{3}^{+}(x,(y_{1},y_{2},0))\left.-\nu\left(\frac{\partial}{\partial y_{1}}\frac{\partial u^{1}}{\partial y_{3}}+\frac{\partial}{\partial y_{2}}\frac{\partial u^{2}}{\partial y_{3}}\right)\right|_{y_{3}=0}(y_{1},y_{2},0)\textrm{d}y_{1}\textrm{d}y_{2}.$
For the last integral we shall perform integration by parts, we deduce that
$\nabla p(x)$ $=\int_{\mathbb{R}_{+}^{3}}K_{3}^{+}(x,y)\Delta p(y)\textrm{d}y+\int_{\mathbb{R}^{2}}K_{3}^{+}(x,(y_{1},y_{2},0))F^{3}(y_{1},y_{2},0)\textrm{d}y_{1}\textrm{d}y_{2}$
$+\nu\int_{\mathbb{R}^{2}}\frac{\partial K_{3}^{+}}{\partial y_{1}}(x,(y_{1},y_{2},0))\frac{\partial u^{1}}{\partial y_{3}}(y_{1},y_{2},0)\textrm{d}y_{1}\textrm{d}y_{2}$
$+\nu\int_{\mathbb{R}^{2}}\frac{\partial K_{3}^{+}}{\partial y_{2}}(x,(y_{1},y_{2},0))\frac{\partial u^{2}}{\partial y_{3}}(y_{1},y_{2},0)\textrm{d}y_{1}\textrm{d}y_{2}.$
Here again those partial derivatives of $K_{3}^{+}$ can be worked out explicitly. ∎
Finally let us mention the integral representation for viscous incompressible flows without constraint.
Lemma 3.
Suppose $p(x,t)$ is the pressure of an incompressible fluid flow in $\mathbb{R}^{d}$ (where $d=2$ or $3$) with viscosity constant $\nu>0$. Then
$\nabla p(x)=\int_{\mathbb{R}^{d}}K_{d}(x,y)\Delta p(y)\textrm{d}y\quad\textrm{for }x\in\mathbb{R}^{d},$ (24)
where $K_{d}(x,y)=\nabla_{x}\Gamma_{d}(x-y)$ is the Biot-Salve singular integral kernel, given by explicitly
$K_{d}(x,y)=\frac{1}{s_{d-1}}\frac{x-y}{|x-y|^{d}}\quad\textrm{ for }x\neq y,$ (25)
and $s_{d-1}$ is the area of the $d-1$ dimensional unit sphere, $s_{1}=2\pi$ and $s_{2}=4\pi$.
2.2 Representing the velocity in terms of fluid particles
After obtaining the pressure gradient $\nabla p$ for any fixed $t$ we may update the velocity field $u(x,t)=(u^{1}(x,t),\cdots,u^{d}(x,t))$ by considering $\nabla p$ as an additional force term applying to the flow. The Navier-Stokes equations may be written as a parabolic equation
$\left(L_{-u}-\frac{\partial}{\partial t}\right)u+g=0\quad\textrm{ in }D,$ (26)
subject to the non-slip condition that $u(x,t)=0$ for $x\in\partial D$ and the initial condition that $u(x,0)=u_{0}(x)$, where $L_{-u}=\nu\Delta-u\cdot\nabla$ and $g=-\nabla p+F$. Hence the velocity $u(x,t)$ may be represented in terms of the distribution of the Brownian fluid particles. Recall that the Brownian fluid particles with velocity $u(x,t)$ are governed by Itô’s stochastic differential equation
$\textrm{d}X_{t}=u(X_{t},t)\textrm{d}t+\sqrt{2v}\textrm{d}B_{t},\quad X_{0}=\eta,$ (27)
where $B$ is $d$-dimensional Brownian motion on a probability space $(\varOmega,\mathcal{F},\mathbb{P})$, and $\eta\in\mathbb{R}^{d}$ is the initial position of Brownian fluid particles.
The particles defined by (27) initiated from $\eta$ is denoted by $X^{\eta}$. Brownian particles defined by (27) form a family of diffusion with the infinitesimal generator $L_{u}=\nu\Delta+u\cdot\nabla$. Let $p_{u}(s,x;t,y)$ be the transition probability density function of Brownian fluid particles, that is, $p_{u}(s,x;t,y)=\mathbb{P}\left[\left.X_{t}=y\right|X_{s}=x\right]$ (for $t>s$) the probability of the fluid particles being at $y$ at time $t$ for those particles initiated from $x$ at time $s$.
Theorem 4.
Suppose $u(x,t)$ is the velocity of an viscous incompressible flow in $\mathbb{R}^{d}$. Then
$u(\xi,t)=\int_{\mathbb{R}^{d}}p_{u}(0,\eta;t,\xi)u_{0}(\eta)\textrm{d}\eta+\int_{0}^{t}\int_{\mathbb{R}^{d}}\mathbb{E}\left[\left.g(X_{s}^{\eta},s)\right|X_{t}^{\eta}=\xi\right]p_{u}(0,\eta;t,\xi)\textrm{d}\eta\textrm{d}s$ (28)
for any $\xi\in\mathbb{R}^{d}$ and $t\geq 0$, where $u_{0}=u(\cdot,0)$ is the initial velocity.
Let us now consider a flow past a flat plate, i.e. an viscous incompressible flow $u(x,t)=(u^{1}(x,t),\cdots,u^{d}(x,t))$ for $x\in\mathbb{R}_{+}^{d}$, subject to the non-slip condition that $u(x,t)=0$ for $x=(x_{1},\cdots,x_{d})$ with $x_{d}=0$.
The velocity $u(x,t)$ is extended to the whole space by reflection, and therefore $u(\bar{x},t)=\overline{u(x,t)}$ for every $x\in\mathbb{R}^{d}$. That is, $u^{i}(\bar{x},t)=u^{i}(x,t)$ ($i=1,\cdots,d-1$) and $u^{d}(\bar{x},t)=-u^{d}(x,t)$. The extension $u(x,t)$ is then divergence-free in distribution on $\mathbb{R}^{d}$ for each $t$, and therefore $L_{u}$ is the $L^{2}$-adjoint of $L_{-u}$. Hence $p_{u}(s,x;t,y)$ is the fundamental solution of the backward parabolic operator $L_{-u}+\frac{\partial}{\partial t}$. According to the integral representation (cf. Theorem 9 below), we have the following representation theorem.
Theorem 5.
For a viscous incompressible flow in $\mathbb{R}_{+}^{d}$, the following integral representation holds:
$u(\xi,t)$ $=\int_{\mathbb{R}_{+}^{d}}\left(p_{u}(0,\eta;t,\xi)-p_{u}(0,\bar{\eta};t,\xi)\right)u_{0}(\eta)\textrm{d}\eta$
$+\int_{0}^{t}\int_{\mathbb{R}_{+}^{d}}\mathbb{E}\left[\left.1_{\{t-s<\zeta(X^{\eta})\}}g(X_{s}^{\eta},s)\right|X_{t}^{\eta}=\xi\right]p_{u}(0,\eta;t,\xi)\textrm{d}\eta\textrm{d}s$ (29)
for $\xi\in\mathbb{R}_{+}^{d}$ and $t>0$, and $u(\bar{\xi},t)=\overline{u(\xi,t)}$ for $\xi\notin\mathbb{R}_{+}^{d}$. Here $\zeta(w)=\inf\{s>0:w(s)\notin\mathbb{R}_{+}^{d}\}$ for every continuous path $w$ in $\mathbb{R}^{d}$.
We therefore have formulated the Navier-Stokes equations in terms of (29, 27, 19, 21). We shall use this formulation to implement large eddy simulations for viscous incompressible flows past a solid boundary.
3 Random LES
We borrow a key idea from LES to carry out numerical experiments based on the functional integral representation (29) for viscous incompressible flows. The principal idea in LES is to facilitate the calculation by ignoring the eddies in small length scales, which are very expensive in computing, via filtering of the Navier–Stokes equations. Such filtering, which can be seen as spatial averaging, effectively removes the information of eddies in small-scale and saves the computing cost, cf. [15]. That is, LES method does not aim to calculate the velocity $u(x,t)$, but only aims to compute the local average $\tilde{u}(x,t)$ at every grid point. To implement this approach, we choose a filter function $\chi$, which is a non-negative function on $\mathbb{R}^{d}$, ideally with a compact support, such that $\int_{\mathbb{R}^{d}}\chi(x)\textrm{d}x=1$. A possible choice of the filter is the Gaussian kernel on $\mathbb{R}^{d}$ with variance $\sigma^{2}>0$:
$\chi(x)=\frac{1}{(2\pi\sigma^{2})^{\frac{d}{2}}}\exp\left(-\frac{|x|^{2}}{2\sigma^{2}}\right)\quad\textrm{ for }x\in\mathbb{R}^{d},$
where $\sigma^{2}>0$ should be small and is determined by the grid size used in the numerical schemes. This filter function has its support on $\mathbb{R}^{d}$, while when $\sigma^{2}>0$ is small enough, $\chi(x-y)$ is concentrated at about $x$.
3.1 Random LES for flows in $\mathbb{R}^{d}$
For a flow in $\mathbb{R}^{d}$, the filtered velocity $\tilde{u}(x,t)$ is defined by
$\tilde{u}(x,t)=\int_{\mathbb{R}^{d}}\chi(x-y)u(y,t)\mathrm{d}y\quad\textrm{ for }x\in\mathbb{R}^{d}.$ (30)
The filtered Navier-Stokes equations become
$\frac{\partial}{\partial t}\tilde{u}+\widetilde{(u\cdot\nabla)u}=\nu\Delta\tilde{u}-\nabla\tilde{p}+\tilde{F}$
and
$\nabla\cdot\tilde{u}=0,$
Because $\widetilde{(u\cdot\nabla)u}$ can not be simplified as $(\tilde{u}\cdot\nabla)\tilde{u}$, so people define sub-grid-scale stress $\tau$ as
$\nabla\cdot\tau=(\tilde{u}\cdot\nabla)\tilde{u}-\widetilde{(u\cdot\nabla)u}$
Therefore the filtered Navier-Stokes equation can be rewritten as
$\partial_{t}\tilde{u}+(\tilde{u}\cdot\nabla)\tilde{u}=\nu\Delta\tilde{u}-\nabla\tilde{p}+\tilde{F}-\nabla\cdot\tau.$
The tensor $\tau$ measures the effect of small eddies on the flow system which should be modelled. The influence of small-scale eddies on the equation of motion is described by some other models (cf. [12], [16] and [13]), but it will not be elaborated in our paper. LES effectively reduces the computational complexity and retains the main physical properties.
Lemma 6.
For a viscous incompressible in $\mathbb{R}^{d}$, the following representation holds
$\tilde{u}(x,t)=\int_{\mathbb{R}^{d}}\mathbb{E}\left[\chi(x-X_{t}^{\eta})u_{0}(\eta)\right]\textrm{d}\eta+\int_{0}^{t}\int_{\mathbb{R}^{d}}\mathbb{E}\left[\chi(x-X_{t}^{\eta})g(X_{s}^{\eta},s)\right]\textrm{d}\eta\textrm{d}s$ (31)
for $x\in\mathbb{R}^{d}$.
Proof.
Using the integral representation (28) we obtain that
$\tilde{u}(x,t)$ $=\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\chi(x-y)p_{u}(0,\eta;t,y)u_{0}(\eta)\textrm{d}y\textrm{d}\eta$
$\;+\int_{0}^{t}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\chi(x-y)\mathbb{E}\left[\left.g(X_{s}^{\eta},s)\right|X_{t}^{\eta}=y\right]p_{u}(0,\eta;t,y)\textrm{d}y\textrm{d}\eta\textrm{d}s$
$=\int_{\mathbb{R}^{d}}\mathbb{E}\left[\chi(x-X_{t}^{\eta})u_{0}(\eta)\right]\textrm{d}\eta+\int_{0}^{t}\int_{\mathbb{R}^{d}}\mathbb{E}\left[\chi(x-X_{t}^{\eta})g(X_{s}^{\eta},s)\right]\textrm{d}\eta\textrm{d}s$
which completes the proof. ∎
Let us now write down the numerical scheme of random LES.
Random LES method
Based on this representation, we are able to propose the following random LES scheme for flows without space constrain as the following.
$\textrm{d}Y_{t}^{\eta}=U(Y_{t}^{\eta},t)\textrm{d}t+\sqrt{2\nu}\textrm{d}B_{t},\quad Y_{0}^{\eta}=\eta$ (32)
for every $\eta\in\mathbb{R}^{d}$, where $B=(B^{1},\cdots,B^{d})$ is a standard Brownian motion on a probability space $(\varOmega,\mathcal{F},\mathbb{P})$;
$U(x,t)=\int_{\mathbb{R}^{d}}\mathbb{E}\left[\chi(x-Y_{t}^{\eta})u_{0}(\eta)\right]\textrm{d}\eta+\int_{0}^{t}\int_{\mathbb{R}^{d}}\mathbb{E}\left[\chi(x-Y_{t}^{\eta})G(Y_{s}^{\eta},s)\right]\textrm{d}\eta\textrm{d}s$ (33)
for $x\in\mathbb{R}^{d}$, where $u_{0}$ is the initial velocity,
$G(x,t)=-\nabla P(x,t)+F(x,t)$ (34)
for every $x\in\mathbb{R}^{d}$ and $t\geq 0$, and $F=(F^{1},F^{2},F^{3})$ is an external force applying to the flow;
$\nabla P(x,t)=\int_{\mathbb{R}^{d}}K_{d}(x,y)\left.\left(\nabla\cdot F-\sum_{i,j=1}^{d}\frac{\partial U^{j}}{\partial y^{i}}\frac{\partial U^{i}}{\partial y^{j}}\right)\right|_{(y,t)}\textrm{d}y$ (35)
for every $x\in\mathbb{R}^{d}$ and $t\geq 0$, where $K_{d}(x,y)=\frac{1}{s_{d-1}}\frac{x-y}{|x-y|^{d}}$ for $x\neq y$, $s_{1}=2\pi$ and $s_{2}=4\pi$.
3.2 Random LES for flows past a wall
For a flow constrained in $\mathbb{R}_{+}^{d}$, i.e. for a flow past a flat plate, the velocity field $u(x,t)=(u^{1}(x,t),\cdots,u^{d}(x,t))$ is extended to the whole space $\mathbb{R}^{d}$ by reflection about $x_{d}=0$ so that the relation $u(\bar{x},t)=\overline{u(x,t)}$ is maintained for all $t$ and $x$. Then the filtered velocity is defined to be
$\tilde{u}(x,t)=\int_{\mathbb{R}^{d}}\chi(x-y)u(y,t)\mathrm{d}y$
for every $x\in\mathbb{R}^{d}$ and $t$. Therefore, by a change of variable,
$\tilde{u}^{i}(x,t)=\int_{\mathbb{R}_{+}^{d}}\chi_{+}(x,y)u^{i}(y,t)\mathrm{d}y,\quad\tilde{u}^{d}(x,t)=\int_{\mathbb{R}_{+}^{d}}\chi_{-}(x,y)u^{d}(y,t)\mathrm{d}y$ (36)
for $i=1,\cdots,d-1$, and any $x$ and $t$, where
$\chi_{\pm}(x,y)=\chi(x-y)\pm\chi(x-\bar{y})$ (37)
for any $x,y\in\mathbb{R}^{d}$. In other words
$\tilde{u}^{i}(x,t)=\int_{\mathbb{R}_{+}^{d}}\chi_{i}(x,y)u^{i}(y,t)\mathrm{d}y$
for all $i=1,\cdots,d$, where $\chi_{i}=\chi_{+}$ for $i=1,\cdots,d-1$ and $\chi_{d}=\chi_{-}$.
Lemma 7.
For a flow in $\mathbb{R}_{+}^{d}$, it holds that
$\tilde{u}^{i}(x,t)$ $=\int_{\mathbb{R}_{+}^{d}}\left(\mathbb{E}\left[\chi_{i}(x,X_{t}^{\eta})1_{\left\{t<\zeta\right\}}\right]-\mathbb{E}\left[\chi_{i}(x,X_{t}^{\bar{\eta}})1_{\left\{t<\zeta\right\}}\right]\right)u_{0}^{i}(\eta)\textrm{d}\eta$
$\;+\int_{0}^{t}\int_{\mathbb{R}_{+}^{d}}\mathbb{E}\left[1_{\left\{t<\zeta,t-s<\zeta\circ\tau_{t}\right\}}\chi_{i}(x,X_{t}^{\eta})g^{i}(X_{s}^{\eta},s)\right]\textrm{d}\eta\textrm{d}s$
for $x\in\mathbb{R}_{+}^{d}$ and $t>0$, $\chi_{i}=\chi_{+}$ for $i=1,\cdots,d-1$ and $\chi_{d}=\chi_{-}$. Here $\tau_{t}$ is the time reverse operator from $t>0$ and $\zeta=\inf\left\{t>0:X_{t}\notin\mathbb{R}_{+}^{d}\right\}$.
Proof.
According to Theorem 5,
$u^{i}(y,t)$ $=\int_{\mathbb{R}_{+}^{d}}\left(p_{u}(0,\eta;t,y)-p_{u}(0,\bar{\eta};t,y)\right)u_{0}^{i}(\eta)\textrm{d}\eta$
$+\int_{0}^{t}\int_{\mathbb{R}_{+}^{d}}\mathbb{E}\left[\left.1_{\{t-s<\zeta(X^{\eta}\circ\tau_{t})\}}g^{i}(X_{s}^{\eta},s)\right|X_{t}^{\eta}=\xi\right]p_{u}(0,\eta;t,y)\textrm{d}\eta\textrm{d}s$ (38)
for all $i=1,\cdots,d$, where $\tau_{t}$ is the time reverse from $t$. By using the Fubini theorem, we have the following integral representation.
$\int_{\mathbb{R}_{+}^{d}}\chi_{i}(x,y)u^{i}(y,t)$ $=\int_{\mathbb{R}_{+}^{d}}\left(\int_{\mathbb{R}_{+}^{d}}\chi_{i}(x,y)p_{u}(0,\eta;t,y)\textrm{d}y-\int_{\mathbb{R}_{+}^{d}}\chi_{i}(x,y)p_{u}(0,\bar{\eta};t,y)\textrm{d}y\right)u_{0}^{i}(\eta)\textrm{d}\eta$
$+\int_{0}^{t}\int_{\mathbb{R}_{+}^{d}}\int_{\mathbb{R}_{+}^{d}}\chi_{i}(x,y)\mathbb{E}\left[\left.1_{\{t-s<\zeta(X^{\eta})\}}g^{i}(X_{s}^{\eta},s)\right|X_{t}^{\eta}=\xi\right]p_{u}(0,\eta;t,y)\textrm{d}y\textrm{d}\eta\textrm{d}s$ (39)
$=\int_{\mathbb{R}_{+}^{d}}\left(\mathbb{E}\left[\chi_{i}(x,X_{t}^{\eta})1_{\left\{t<\zeta\right\}}\right]-\mathbb{E}\left[\chi_{i}(x,X_{t}^{\bar{\eta}})1_{\left\{t<\zeta\right\}}\right]\right)u_{0}^{i}(\eta)\textrm{d}\eta$ (40)
$+\int_{0}^{t}\int_{\mathbb{R}_{+}^{d}}\mathbb{E}\left[1_{\left\{t<\zeta\right\}}1_{\{t-s<\zeta\circ\tau_{t}\}}\chi_{i}(x,X_{t}^{\eta})g^{i}(X_{s}^{\eta},s)\right]\textrm{d}\eta\textrm{d}s$ (41)
which completes the proof. ∎
This representation suggests the following approximation.
Random LES method for flows past walls
Run a diffusion with vector field $U(x,t)$:
$\textrm{d}Y_{t}^{\eta}=U(Y_{t}^{\eta},t)\textrm{d}t+\sqrt{2\nu}\textrm{d}B_{t},\quad Y_{0}^{\eta}=\eta$ (42)
for every $\eta\in\mathbb{R}^{d}$, where $B=(B^{1},\cdots,B^{d})$ is a standard Brownian motion on a probability space $(\varOmega,\mathcal{F},\mathbb{P})$;
$U^{i}(x,t)$ $=\int_{\mathbb{R}_{+}^{d}}\left(\mathbb{E}\left[\chi_{i}(x,Y_{t}^{\eta})1_{\left\{t<\zeta\right\}}\right]-\mathbb{E}\left[\chi_{i}(x,Y_{t}^{\bar{\eta}})1_{\left\{t<\zeta\right\}}\right]\right)u_{0}^{i}(\eta)\textrm{d}\eta$
$\;+\int_{0}^{t}\int_{\mathbb{R}_{+}^{d}}\mathbb{E}\left[1_{\left\{t<\zeta,t-s<\zeta\circ\tau_{t}\right\}}\chi_{i}(x,Y_{t}^{\eta})G^{i}(Y_{s}^{\eta},s)\right]\textrm{d}\eta\textrm{d}s$ (43)
for $x\in\mathbb{R}_{+}^{d}$, $i=1,\cdots,d$, and
$U(x,t)=\overline{U(\bar{x},t)}\quad\textrm{ for }x\in\mathbb{R}_{-}^{d},\quad U(x,t)=0\quad\textrm{ if }x_{d}=0,$ (44)
where $u_{0}$ is the initial velocity;
$G(x,t)=-\nabla P(x,t)+F(x,t)\quad\textrm{ for }x\in\mathbb{R}_{+}^{d},t\geq 0,$ (45)
where $F=(F^{1},F^{2},F^{3})$ is an external force applying to the flow, and
(1) if $d=2$,
$\nabla P(x,t)$ $=\int_{\mathbb{R}_{+}^{2}}K_{2}^{+}(x,y)\left.\left(\nabla\cdot F-\sum_{i,j=1}^{2}\frac{\partial U^{j}}{\partial y^{i}}\frac{\partial U^{i}}{\partial y^{j}}\right)\right|_{(y,t)}\textrm{d}y$
$\;+\int_{\mathbb{R}}K_{2}^{+}(x,(y_{1},0))F^{2}(y_{1},0)\textrm{d}y_{1}$
$\;+\nu\int_{\mathbb{R}}\frac{\partial K_{2}^{+}}{\partial y_{1}}(x,(y_{1},0))\frac{\partial U^{1}}{\partial y_{2}}(y_{1},0)\textrm{d}y_{1},$ (46)
(2) if $d=3$,
$\nabla P(x,t)$ $=\int_{\mathbb{R}_{+}^{3}}K_{3}^{+}(x,y)\left.\left(\nabla\cdot F-\sum_{i,j=1}^{3}\frac{\partial U^{j}}{\partial y^{i}}\frac{\partial U^{i}}{\partial y^{j}}\right)\right|_{(y,t)}\textrm{d}y$
$\;+\int_{\mathbb{R}^{2}}K_{3}^{+}(x,(y_{1},y_{2},0))F^{3}(y_{1},y_{2},0)\textrm{d}y_{1}\textrm{d}y_{2}$
$\;+\nu\int_{\mathbb{R}^{2}}\frac{\partial K_{3}^{+}}{\partial y_{1}}(x,(y_{1},y_{2},0))\frac{\partial U^{1}}{\partial y_{3}}(y_{1},y_{2},0)\textrm{d}y_{1}\textrm{d}y_{2}$
$\;+\nu\int_{\mathbb{R}^{2}}\frac{\partial K_{3}^{+}}{\partial y_{2}}(x,(y_{1},y_{2},0))\frac{\partial U^{2}}{\partial y_{3}}(y_{1},y_{2},0)\textrm{d}y_{1}\textrm{d}y_{2}$ (47)
for $x\in\mathbb{R}_{+}^{d}$ and $t\geq 0$.
4 Numerical schemes and experiments
4.1 Numerical schemes and variables settings
In this section, we present comprehensive numerical experiments based on the scheme defined in Section 3. We are able to present the numerical scheme based on above formulas, and we use the Euler method to discrete the integral. The initial velocity $U(x,0)$ and external force $F$ is given, so the $\nabla\cdot F$ can be calculated directly.
Set mesh size $s>0$, time step $\delta>0$ and kinematic viscosity $\nu>0$. For $i_{1},i_{2},i_{3}\in\mathbb{Z}$, denote $y^{i_{1,}i_{2},i_{3}}=(i_{1},i_{2},i_{3})s$, $U^{i_{1},i_{2},i_{3}}=U(y^{i_{1,}i_{2},i_{3}},0)$. In numerical scheme, we drop the expectation and use one-copy of Brownian particles. We discrete the stochastic differential equation (42) following Euler scheme: for $t_{i}=i\delta,i=0,1,2,\cdots$.
$Y_{t_{k}}^{i_{1},i_{2},i_{3}}=Y_{t_{k-1}}^{i_{1},i_{2},i_{3}}+\delta U(Y_{t_{k-1}}^{i_{1},i_{2},i_{3}},t_{k-1})+\sqrt{2\nu}(B_{t_{k}}-B_{t_{k-1}}),\quad Y_{0}^{i_{1},i_{2},i_{3}}=y^{i_{1,}i_{2},i_{3}}$
where $Y_{t_{k}}^{i_{1},i_{2},i_{3}}$ denotes $Y_{t_{k}}^{x^{i_{1},i_{2},i_{3}}}$ for simplicity. Then integral representations in (43) and (47) can be discretized as follows:
$U(x,t_{k})$ $=\sum_{i_{1},i_{2}\in\mathbb{R},i_{3}>0}s^{3}\left[1_{\mathbb{R}_{+}^{3}}\left(Y_{t_{l}}^{i_{1},i_{2},i_{3}}\right)\chi_{i}(x-Y_{t_{k}}^{i_{1},i_{2},i_{3}})-1_{\mathbb{R}_{+}^{3}}\left(Y_{t_{l}}^{i_{1},i_{2},i_{3}}\right)\chi_{i}(x-Y_{t_{k}}^{i_{1},i_{2},-i_{3}})\right]U_{0}^{i_{1},i_{2},i_{3}}$
$+\sum_{i_{1},i_{2}\in\mathbb{R},i_{3}>0}\sum_{j=1}^{k}s^{3}\delta\prod_{l=j}^{k}1_{\mathbb{R}_{+}^{3}}\left(Y_{t_{l}}^{i_{1},i_{2},i_{3}}\right)\chi_{i}(x-Y_{t_{k}}^{i_{1},i_{2},i_{3}})G(X_{t_{j-1}}^{i_{1},i_{2},i_{3}},t_{j-1})$ (48)
for $x\in\mathbb{R}_{+}^{d}$, and
$U(x,t)=\overline{U(\bar{x},t)}\quad\textrm{ for }x\in\mathbb{R}_{-}^{d},\quad U(x,t)=0\quad\textrm{ for }x_{3}=0,$ (49)
where $G(x,t_{k})=-\nabla P(x,t_{k})+F(x,t_{k})$,
$\nabla P(x,t_{k})$ $=\sum_{i_{1},i_{2}\in\mathbb{R},i_{3}>0}K_{3}^{+}(x,(i_{1},i_{2},i_{3}))\left(\nabla\cdot F((i_{1},i_{2},i_{3}),t_{k})-\sum_{i,j=1}^{3}\frac{\partial U^{j}}{\partial x^{i}}((i_{1},i_{2},i_{3}),t_{k})\frac{\partial U^{i}}{\partial x^{j}}((i_{1},i_{2},i_{3}),t_{k})\right)$
$\;+\sum_{i_{1},i_{2}\in\mathbb{R}}s^{2}K_{3}^{+}(x,(i_{1},i_{2},0))F^{3}(i_{1},i_{2},0)$
$\;+\nu\sum_{i_{1},i_{2}\in\mathbb{R}}s^{2}\frac{\partial K_{3}^{+}}{\partial y_{1}}(x,(i_{1},i_{2},0))\frac{\partial U^{1}}{\partial x_{3}}((i_{1},i_{2},0),t_{k})$
$\;+\nu\sum_{i_{1},i_{2}\in\mathbb{R}}s^{2}\frac{\partial K_{3}^{+}}{\partial y_{2}}(x,(i_{1},i_{2},0))\frac{\partial U^{2}}{\partial x_{3}}((i_{1},i_{2},0),t_{k})$ (50)
for $x=(x_{1},x_{2},x_{3})$ where $x_{3}>0$, the $\frac{\partial U^{j}}{\partial x^{i}}$ can be explicitly represented and calculated directly.
To validate the correctness of the theory and numerical methods, we conducted extensive numerical experiments. For comprehensiveness, the following experiments included scenarios with both laminar flows and turbulence, in two-dimensional and three-dimensional settings, as well as in unbounded domains and with wall bounded space. From the perspective of numerical simulation, turbulence represents a more challenging flow regime to model. By comparing the results of laminar flow experiments with those of turbulent flow experiments, we can more clearly discern the differences between these two flow states. The results demonstrate that random LES method can accurately and efficiently simulate fluid systems.
Now we introduce some fundamental settings. Reynolds number $Re$ is defined by $Re=\frac{U_{0}L}{\nu}$, where $U_{0}$ is the main stream velocity, $\nu$ is kinematic viscosity and $L$ is the length scale. In unbounded experiments, the observation region is set as a square, meaning that the horizontal length scale ($L_{h}$) and vertical length scale ($L_{v}$) of the observation area are equal. For the space with solid wall boundary, a smaller observation domain and a more refined grid discretization are employed in the vertical direction to enhance the resolution of boundary phenomena. The mesh size $s\thicksim L\sqrt{\frac{1}{Re}}$, we set the vertical mesh size as $s_{v}$ and the horizontal mesh size as $s_{h}$ and the external force as $F$.
In unbounded experiments, it is essential to introduce an initial vorticity into the fluid system to observe its subsequent evolution. This necessity arises because, in the absence of initial vorticity, an unbounded fluid system subjected solely to linear forces will not exhibit vortex dynamics. In contrast, in experiments involving wall-bounded systems, vorticity can naturally emerge due to boundary pressure effects, particularly near the walls. Furthermore, to facilitate the observation of vorticity formation, it is advisable to apply specific nonlinear forces concurrently. Therefore, in the subsequent experiments, we adopt the following setup: (1) for unbounded systems, an initial rotational velocity field is imposed along with linear forces; (2) for wall-bounded systems, an initial uniform velocity field (devoid of vorticity) is prescribed, accompanied by the application of nonlinear forces.
The experiments conducted without boundary conditions (4.2.1, 4.2.3, 4.3.1, 4.3.3) serve as a simplified "toy model" for preliminary exploration. Moving forward, our focus will shift to more realistic models that incorporate boundary conditions (4.2.2, 4.2.4, 4.3.2, 4.3.4). Through a comprehensive set of experiments, we aim to demonstrate the effectiveness of our approach.
4.2 Experiments for laminar flows
A laminar flow refers to a viscous flow characterized by a low Reynolds number. It is a smooth and orderly fluid motion in which the fluid moves in parallel layers without mixing between them. The mathematical description of laminar flow is relatively straightforward. Consequently, compared to turbulent flow, laminar flow requires less computational complexity and lower resource demands for numerical simulations.
4.2.1 2-dimensional unbounded laminar flows
In this experiment, we set $\nu=0.3$, $Re=1000$, length scale $L=3\pi$, so that $U_{0}=\frac{\nu}{L}Re=31.83$. The mesh size $s=\frac{3\pi}{50}\thicksim L\sqrt{\frac{1}{Re}}$ and the time step $\delta=0.001$. The numerical experiment is demonstrated at times $t_{1}=0.03,t_{2}=0.06,t_{3}=0.09$. We set the initial velocity to be of the form $U(x,0)=(U_{0}\sin(2x_{2}),U_{0}\cos(2x_{1}))$, and set force $F=(10,-9.81)$. The velocity field and the vorticity field are shown in Figure 1.
Figure 1: Velocity fields (left) and vorticity (right) of unbounded laminar flows on $\mathbb{R}^{2}$
Since no external boundary conditions or nonlinear forces were applied in this experiment, an initial rotating vector field is necessary to observe the system’s evolution. As illustrated in Figure 1, the velocity field transitions from an initially rotating pattern to a more irregular state, eventually entering a chaotic regime. Correspondingly, the vorticity exhibits fluctuations that align with the expected behavior predicted by our experimental setup and hypothesis.
4.2.2 2-dimensional wall-bounded laminar flows
Now we consider the bounded case. In this part, $\nu=0.3$, $Re=1000$, length scale $L_{h}=3\pi,L_{v}=0.2\pi$, so that $U_{0}=\frac{\nu}{L_{h}}Re=31.83$. The mesh size $s_{h}=\frac{3\pi}{50},s_{v}=\frac{0.2\pi}{50}$. and the time step $\delta=0.001$. The numerical experiment is demonstrated at times $t=0.03,t=0.06,t=0.09$. We set the initial velocity to be of the form $U(x,0)=(U_{0},0)$, and set force $F=(10e^{(-\frac{(x_{1}^{2}+x_{2}^{2})}{2s_{h}s_{v}})},-9.81)$. The velocity field and the vorticity field are shown in Figure 1.
Figure 2: Velocity fields (left) and vorticity (right) of wall-bounded laminar flows on $\mathbb{R}_{+}^{2}$
As shown in Figure 2, the fluid system initially remains in a relatively stable state. Over time, a small number of vortices begin to emerge at the boundary and gradually propagate into the interior of the domain. This phenomenon is primarily driven by shear forces, specifically the combined effects of boundary pressure and nonlinear forces. Notably, compared to the subsequent turbulent flow experiments (Figure 12), the fluid dynamics in the laminar flow experiments exhibit significantly lower intensity, with fewer vortices generated. This observation is consistent with established theoretical principles.
4.2.3 3-dimensional unbounded laminar flows
Similar to 2-dimensional numerical experiments, we set $\nu=0.3$, $Re=1000$, length scale $L=3\pi$, so that $U_{0}=\frac{\nu}{L}Re=31.83$. The mesh size $s=\frac{3\pi}{50}\thicksim L\sqrt{\frac{1}{Re}}$. and the time step $\delta=0.001$. The numerical experiment is demonstrated at times $t=0.03,t=0.06,t=0.09$. We set the initial velocity to be of the form $U(x,0)=(U_{0}\sin(2x_{2}),U_{0}\cos(2x_{1}),0)$, and set force $F=(10,10,-9.81)$. The velocity field and the vorticity field are shown in Figure 3.
Figure 3: Velocity fields (left) and vorticity (right) of unbounded laminar flows on $\mathbb{R}^{3}$
Similar to the unbounded two-dimensional experiment, the initial rotating velocity field undergoes a transition from its initial formation to gradual dissipation. As shown in Figure 3, the initial conditions generate a vortex-like flow pattern, while the vorticity exhibits increasingly chaotic behavior. To enhance the visualization of the flow dynamics, sectional views of the velocity fields are presented at the planes $x=C$, $y=C$, and $z=C$ in Figures 4, 5, and 6, respectively, with $C=0.1\pi$ chosen for illustration.
Figure 4: Section Velocity fields of incompressible viscous flows on plane on $x=C$
Figure 5: Section Velocity fields of incompressible viscous flows on plane on $y=C$
Figure 6: Section Velocity fields of incompressible viscous flows on plane on $z=C$
The sectional velocity fields are influenced by the initial conditions specified in the setup. Based on the initial velocity field configuration, the flow on $z=C$ plane exhibits more complex behavior, as illustrated in Figure 6. From the sectional views, it is observed that as time progresses, the fluid predominantly moves in the direction of the applied force, consistent with the expected physical behavior.
4.2.4 3-dimensional wall-bounded laminar flows
In this experiment, $\nu=0.3$, $Re=1000$, $L_{h}=3\pi,L_{v}=0.2\pi$, so that $U_{0}=\frac{\nu}{L}Re=31.83$. The horizontal mesh size $s_{h}=\frac{3\pi}{50}$, vertical mesh size $s_{v}=\frac{0.2\pi}{50}$ , and the time step $\delta=0.001$. The numerical experiment is demonstrated at times $t=0.03,t=0.06,t=0.09$. We set the initial velocity to be of the form $U(x,0)=(U_{0},U_{0},0)$, and set force $F=(100e^{(-\frac{(x_{1}^{2}+x_{2}^{2}+x_{3}^{2})}{2s_{h}^{2}s_{v}})},100e^{(-\frac{(x_{1}^{2}+x_{2}^{2}+x_{3}^{2})}{2s_{h}^{2}s_{v}})},-9.81)$. The velocity field and the vorticity field are shown in Figure 7.
Figure 7: Velocity fields (left) and vorticity (right) of wall-bounded laminar flows on $\mathbb{R}_{+}^{3}$
In contrast to the unbounded case, the presence of gravity and boundary shear forces induces fluctuations in the flow near the boundary, leading to the formation of vorticity. As a result, the absolute value of the boundary vorticity increases, as shown in Figure 7. We find that the formation of the vortex is not so complete, which is because the changes in laminar flows are not so intense.
Figure 8: Section Velocity fields of incompressible viscous flows on plane on $x=C$
Figure 9: Section Velocity fields of incompressible viscous flows on plane on $y=C$
Figure 10: Section Velocity fields of incompressible viscous flows on plane on $z=C$
The three sectional figures illustrate that on the tangent plane adjacent to the boundary, fluctuations and variations occur to some extent. However, these changes are not significant due to the nature of the laminar flow. Moreover, the fluid flow on the $z=C$ plane largely maintains its initial state.
4.3 Experiments for turbulent flows
4.3.1 2-dimensional unbounded turbulent flows
In this subsection, we present a numerical experiment involving a 2-dimensional turbulent flow in $\mathbb{R}^{2}$. For this experiment, we set the kinematic viscosity $\nu=0.3$ and the Reynolds number $Re=5500$. The horizontal and vertical length scales ($L_{h}$ and $L_{v}$) are both set to $L_{h}=L_{v}=3\pi$, From the relationship between velocity and the Reynolds number, we calculate the initial mean stream velocity as $U_{0}=\frac{\nu}{L}Re=135.15$. The computational grid is defined with a horizontal mesh size $s_{h}=\frac{3\pi}{50}\thicksim L\sqrt{\frac{1}{Re}}$ and a vertical mesh size $s_{v}=\frac{3\pi}{50}$. The time step is set to $\delta=0.0001$. The numerical results are presented at three time instances: $t=0.01,t=0.02,t=0.03$. The initial velocity field is specified as $U(x,0)=(U_{0}sin(2x_{2}),U_{0}cos(2x_{1}))$ and the external force is given by $F=(10,-9.81)$. The velocity fields are shown in Figure 11 (left), and the corresponding vorticity is displayed in Figure 11 (right).
Figure 11: Velocity fields (left) and vorticity (right) of unbounded turbulent flows on $\mathbb{R}^{2}$
Since no external boundary conditions or nonlinear forces were applied in this experiment, an initial rotating vector field is required to observe the system’s evolution. This rationale underpins the chosen form for the initial velocity field. As shown in Figure 11, the velocity field evolves from an initial rotating pattern into a more irregular state, eventually transitioning into a chaotic regime. Correspondingly, the vorticity exhibits fluctuations, consistent with the expected behavior based on our experimental setup and hypothesis.
4.3.2 2-dimensional wall-bounded turbulent flows
In this subsection, we present a numerical experiment simulating a two-dimensional turbulent flow on $\mathbb{R}_{+}^{2}$, where the flow passes over a solid wall at a high Reynolds number. To better capture the boundary effects, we use a smaller observation domain and finer spatial grid resolution in the vertical direction. The parameters for the experiment are as follows: $\nu=0.3$, $Re=5500$, horizontal length scale $L_{h}=3\pi$, vertical length scale $L_{v}=0.2\pi$ and the initial main stream velocity $U_{0}=135.15$. The horizontal mesh size $s_{h}=\frac{3\pi}{50}$, vertical mesh size $s_{v}=\frac{0.2\pi}{50}$ and the time step $\delta=0.0001$. The numerical results are demonstrated at three time instances: $t=0.01,t=0.02,t=0.03$. The initial velocity field is defined as $U(x,0)=(U_{0},0)$ while the external force is given by $F=(100e^{(-\frac{(x_{1}^{2}+x_{2}^{2})}{2s_{h}s_{v}})},-9.81)$. The velocity and vorticity are shown in Figure 12.
Figure 12: Velocity fields (left) and vorticity (right) of wall-bounded turbulent flows on $\mathbb{R}_{+}^{2}$
As shown in Figure 12, the fluid system initially remains in a relatively stable state. Over time, vortices begin to form at the boundary and gradually propagate into the interior of the domain. This phenomenon is primarily attributed to shear forces, specifically the combined influence of boundary pressure and nonlinear forces. In contrast to the boundary-free case, the presence of a solid boundary inherently facilitates vortex generation at the wall. Additionally, the vorticity near the boundary exhibits significantly higher values, which aligns with physical observations in real-world fluid flows.
4.3.3 3-dimensional unbounded turbulent flows
Numerical experiment simulating a three-dimensional turbulent flow at a high Reynolds number is shown in this subsection. In this experiment, the kinematic viscosity is set to $\nu=0.3$ and the Reynolds number is $Re=5500$. Similar to the two-dimensional simulations, the characteristic length scales in both directions are $L_{h}=L_{v}=3\pi$, which gives an initial mean stream velocity of $U_{0}=135.15$. The mesh size $s=\frac{3\pi}{50}$, and the time step $\delta=0.0001$. Numerical results are presented at the following time instances: $t=0.01,t=0.02,t=0.03$. The initial velocity field is defined as $U(x,0)=(U_{0}\sin(2x_{2}),U_{0}\cos(2x_{1}),0)$, and the external force is given by $F=(100,100,-9.81)$. Figure 17 shows the results of the experiment.
Figure 13: Velocity fields (left) and vorticity (right) of unbounded turbulent flows on $\mathbb{R}^{3}$
Similar to the unbounded two-dimensional experiment, in the three-dimensional case, we must either set the initial velocity field as a rotating vector field or add the nonlinear force to observe fluid dynamics. In this experiment, we choose the former approach. As shown in Figure 13, the initial conditions result in a vortex-like flow pattern, while the vorticity exhibits more chaotic behavior. Additionally, as time progresses, the fluid generally flows in the direction of the applied force, which aligns with expected physical behavior. To visualize the flow more clearly, we provide sectional views of the velocity fields at planes $x=C$, plane $y=C$, and the plane $z=C$ in Figure 14, Figure 15, Figure 16 respectively (we choose $C=0.1\pi$).
Figure 14: Section Velocity fields of incompressible viscous flows on plane on $x=C$
Figure 15: Section Velocity fields of incompressible viscous flows on plane on $y=C$
Figure 16: Section Velocity fields of incompressible viscous flows on plane on $z=C$
As shown in Figures 14, 15, and 16, the fluid system gradually becomes chaotic and disordered on the $x=C$ and $y=C$ planes, as only linear forces and trivial initial conditions are applied in these two directions. In contrast, on the $z=C$ plane, the vortex phenomenon is more pronounced due to the initial rotational conditions specified in this plane.
4.3.4 3-dimensional wall-bounded turbulent flows
In this subsection, we present a numerical experiment simulating a three-dimensional turbulent flow in $\mathbb{R}_{+}^{3}$, where the flow passes over a solid wall at a high Reynolds number. To better capture the boundary effects, we use a finer spatial grid resolution in the vertical direction. The experiment is conducted with the following parameters: the kinematic viscosity $\nu=0.3$, $Re=5500$, horizontal length scale $L_{h}=3\pi$, vertical length scale $L_{v}=0.2\pi$, initial main stream velocity $U_{0}=135.15$. The horizontal mesh size $s_{h}=\frac{3\pi}{50}$, vertical mesh size $s_{v}=\frac{0.2\pi}{50}$and the time step $\delta=0.0001$. The numerical experiment results are demonstrated at times $t=0.01,t=0.02,t=0.03$. The initial velocity field is given by $U(x,0)=(U_{0},U_{0},0)$, set force is given by $F=(100e^{(-\frac{(x_{1}^{2}+x_{2}^{2}+x_{3}^{2})}{2s_{h}^{2}s_{v}})},100e^{(-\frac{(x_{1}^{2}+x_{2}^{2}+x_{3}^{2})}{2s_{h}^{2}s_{v}})},-9.81)$. The velocity and vorticity are shown in Figure 17.
Figure 17: Velocity fields (left) and vorticity (right) of wall-bounded turbulent flows on $\mathbb{R}_{+}^{3}$
Figure 17 shows that, in the 3D turbulent flow, the velocity field evolves over time, with vortices gradually forming at the boundary $z=0$. The vorticity field exhibits corresponding changes, which aligns with both theoretical predictions and real-world observations. To better visualize the flow, we provide cross-sectional views of the velocity field at the planes $x=C$, $y=C$, and $z=C$ in Figures 18, 19, and 20, respectively, where $C=0.1\pi$.
Figure 18: Section Velocity fields of incompressible viscous flows on plane $x=C$
Figure 19: Section Velocity fields of incompressible viscous flows on plane $y=C$
Figure 20: Section Velocity fields of incompressible viscous flows on plane $z=C$
In Figures 18 and 19, the application of external forces leads to the formation of eddies at the boundary, which then propagate throughout the domain as the fluid flows across the wall surface. In contrast, on the $z=C$ plane (Figure 20), where no initial nonlinear forces are applied in this direction, only minor changes in the flow are observed, consistent with the initial conditions we set. These experimental results are in agreement with both theoretical predictions and empirical observations, thereby validating our model and numerical scheme.
In general, compared to laminar flow, turbulent flow exhibits a non-uniform velocity distribution with more pronounced gradient variations. Furthermore, turbulent flow is characterized by significant energy dissipation and inherent instability. It generates a greater number of vortices near the wall boundary, which tend to be more fully developed in structure. In contrast, laminar flow produces fewer vortices, which may not fully develop or maintain a complete structure. The real-world phenomena described above are effectively replicated in our experiments, validating the effectiveness and reliability of our methodology.
5 Appendix: Fluid flows represented by Brownian fluid particles
In this section we establish the technical facts about divergence-free vector fields, and their Taylor’s diffusion, which are used in the paper.
The following convention will be applied throughout the paper unless otherwise specified. Suppose $b(x,t)$ be a time-dependent vector field on $\mathbb{R}_{+}^{d}$ such that $\nabla\cdot b=0$ and $b(x,t)=0$ for $x=(x_{1},\cdots,x_{d})$ with $x_{d}=0$, then $b(x,t)$ is extended to the whole space $\mathbb{R}^{d}$ by reflection: $b^{i}(\bar{x},t)=b^{i}(x,t)$ for $i=1,\cdots,d-1$ and $b^{d}(\bar{x},t)=-b^{d}(x,t)$ for any $x\in\mathbb{R}^{d}$ and $t\geq 0$, where $x\mapsto\bar{x}=(x_{1},\cdots,x_{d-1},-x_{d})$ is the reflection about $x_{d}=0$. The divergence-free feature of $u$ is retained though in weak sense, that is, $\nabla\cdot b=0$ in distribution on $\mathbb{R}^{d}$, which implies that the $L^{2}$-adjoint of the (forward) heat operator $L_{b}-\frac{\partial}{\partial t}$ coincides with the (backward) heat operator $L_{-b}+\frac{\partial}{\partial t}$. Here, given a time-dependent vector field $b(x,t)$, $L_{b}=\nu\Delta+b\cdot\nabla$ denotes the time-dependent elliptic operator of second-order in $\mathbb{R}^{d}$. Following an idea of Taylor [19] that determining a vector field $b(x,t)$ is equivalent to the description of ‘imaginary’ fluid particles with the velocity $b(x,t)$, we may consider the diffusions defined by Itô’s stochastic differential equation
$\textrm{d}X=b(X,t)\textrm{d}t+\sqrt{2\nu}\textrm{d}B,$ (51)
where $B$ is a Brownian motion on some probability space. The weak solution (cf. [18] and [8]) of (51) defines the diffusion with the infinitesimal generator $L_{b}$, cf. [6]. Let $p_{b}(s,x;t,y)$ and $p_{b}^{+}(s,x;t,y)$ (for $t>s\geq 0$) be the transition probability density function of the $L_{b}$-diffusion and, respectively, of the $L_{b}$-diffusion stopped on leaving the region $\mathbb{R}_{+}^{d}$. Formally $p_{b}(s,x;t,y)$ is the probability that the diffusion $X_{t}$ hits $y$ given that $X_{s}=x$, and similarly $p_{b}^{+}(s,x;t,y)$ the probability that $X_{t}$ hits $y$ before leaving the domain $\mathbb{R}_{+}^{d}$ given that $X_{s}=x$. Then it holds that
$p_{b}^{+}(s,x;t,y)=p_{b}(s,x;t,y)-p_{b}(s,x;t,\bar{y})\quad\forall x,y\in\mathbb{R}_{+}^{d}.$ (52)
Given $\eta\in\mathbb{R}^{d}$, the distribution of $X$, a weak solution to (51) that $X_{0}=\eta$, is denoted by $\mathbb{P}^{\eta}$, which is a probability measure on the space $C$([0,)$,\mathbb{R}^{d})$ ) of continuous paths. Consider for every $\xi$ , Itô’s stochastic differential equation
$\textrm{d}X_{t}=b(X_{t},t)\textrm{d}t+\sqrt{2\nu}\textrm{d}B_{t},\quad X_{0}=\xi,$ (53)
where $B_{t}$ is $d$-dimensional Brownian motion. In addition to Itô’s SDE formulation, we shall use the weak solution formulation as well, cf. [18]. Let $\varOmega=C([0,\infty),\mathbb{R}^{d})$ be the continuous path space in $\mathbb{R}^{d}$, and $X=(X_{t})_{t\geq 0}$ denote the coordinate process, that is, for each $t\geq 0$, $X_{t}:\varOmega\mapsto\mathbb{R}^{d}$, which sends each path $\psi\in\varOmega$ to $\psi(t)$. For simplicity, $X_{t}$ shall be written as $\psi(t)$ if no confusion may arise from the context. Let $\mathcal{F}_{t}^{0}=\sigma\{X_{s}:s\leq t\}$ be the smallest $\sigma$-algebra on $\varOmega$ such that $X_{s}$ are measurable for all $s\leq t$, and $\mathcal{F}^{0}=\sigma\{X_{s}:s<\infty\}$. Then $\mathcal{F}^{0}=\mathcal{B}(\varOmega)$ the Borel $\sigma$-algebra on $\varOmega$ generated by the uniform convergence over any bounded subset of $[0,\infty)$.
For simplicity let us assume that $b(x,t)$ is bounded and jointly Borel measurable. According to [18], for each $\xi\in\mathbb{R}^{d}$ and each $\tau\geq 0$, there is a unique probability measure $\mathbb{P}_{b}^{\xi,\tau}$ on $(\varOmega,\mathcal{F}^{0})$ such that $\mathbb{P}_{b}^{\xi,\tau}\left[X_{s}=\xi\textrm{ for }s\leq\tau\right]=1$, and
$M_{t}^{[f]}=f(X_{t},t)-f(X_{\tau},\tau)-\int_{\tau}^{t}(L_{b}f)(X_{s},s)\textrm{d}s$
is a martingale (for $t\geq\tau$ and under the probability $\mathbb{P}_{b}^{\xi,\tau}$) for every $f\in C_{b}^{2,1}(\mathbb{R}^{d}\times[0,\infty))$. For simplicity $\mathbb{P}_{b}^{\xi,0}$ is denoted by $\mathbb{P}_{b}^{\xi}$. The probability measure $\mathbb{P}_{b}^{\xi,\tau}$ is called the $L_{b}$-diffusion, or Taylor’s diffusion, started from $\xi$ at instance $\tau$. $L_{b}$ is also called the infinitesimal generator of the diffusion $\mathbb{P}_{b}^{\xi,\tau}$. Since $L_{b}$ is uniformly elliptic, the transition probability $P_{b}(\tau,\xi;t,\textrm{d}x)$ of the $L_{b}$-diffusion, which is the distribution of $X_{t}$ under $\mathbb{P}_{b}^{\xi,\tau}$, where $t>\tau\geq 0$, has a positive and Hölder’s continuous density denoted by $p_{b}(\tau,\xi;t,x)$ (for $t>\tau\geq 0$, $\xi,x\in\mathbb{R}^{d}$). That is, $p_{b}(\tau,\xi;t,x)\textrm{d}x=\mathbb{P}_{b}^{\xi,\tau}\left[X_{t}\in\textrm{d}x\right]$.
Let $T>0$ be given. The conditional distribution $\mathbb{P}_{b}^{\xi}\left[\left.\cdot\right|X_{T}=\eta\right]$ can be constructed as the following. Let
$q_{b}(s,x;t,y)=\frac{p_{b}(s,x;t,y)p_{b}(t,y;T,\eta)}{p_{b}(s,x;T,\eta)}$
for $0\leq s<t<T$ and $x,y\in\mathbb{R}^{d}$, which is a transition probability density function. The diffusion associated with the transition probability density function $q_{b}$ started from $\xi$ at $\tau=0$ shall be denoted by $\mathbb{P}^{\xi,0\rightarrow\eta,T}$, which shall be consider as the probability measure $\varOmega=C([0,T],\mathbb{R}^{d})$.
Let $b^{T}(x,t)=b(x,(T-t)^{+})$ and $\tau_{T}$ be the time reversal operation on $\varOmega$, that is, $\psi\circ\tau_{T}(t)=\psi(T-t)$ for $t\in[0,T]$. It can be verified that
$p_{-b^{T}}(s,x;t,y)=p_{b}(T-t,y;T-s,x)$ (54)
for $0\leq s<t\leq T$, and $x,y\in\mathbb{R}^{d}$. The duality of conditional diffusion laws has been established in [14].
Lemma 8.
Suppose $b(x,t)$ is divergence-free, that is, $\nabla\cdot b=0$ (for every $t$) in the sense of distribution on $\mathbb{R}^{d}$. Then
$\mathbb{P}_{b}^{\xi,0\rightarrow\eta,T}=\mathbb{P}_{-b^{T}}^{\eta,0\rightarrow\xi,T}\circ\tau_{T}$
for every $\xi,\eta\in\mathbb{R}^{d}$.
From now on, we assume that $b(x,t)$ for $x\in\mathbb{R}_{+}^{d}$ and $t\geq 0$ be a time-dependent vector field on $D$, satisfying that $b(x,t)=0$ for $x\in\partial\mathbb{R}_{+}^{d}$ (the non-slip condition). Suppose that $b(x,t)$ is differentiable up to the boundary $\partial\mathbb{R}_{+}^{d}$ and is bounded. For each $t$, the vector field $b(x,t)$ is extended to the whole space $\mathbb{R}^{d}$ via the reflection about the hyperspace $x_{d}=0$, so that
$b^{i}(x,t)=b^{i}(\bar{x},t)\quad\textrm{ for }i=1,\ldots,d-1\textrm{ and }b^{d}(x,t)=-b^{d}(\bar{x},t).$ (55)
Then $\overline{b(x,t)}=b(\bar{x},t)$ for all $x\in\mathbb{R}^{d}$ and $t\geq 0$. We also assume that $b(x,t)$ is divergence-free on $\mathbb{R}_{+}^{d}$. Then $\nabla\cdot b=0$ in the sense of distribution on $\mathbb{R}^{d}$.
Theorem 9.
Let $\varPsi$ be a solution to the parabolic equation
$\left(L_{-b}-\frac{\partial}{\partial t}\right)\varPsi+g=0\quad\textrm{ in }\mathbb{R}_{+}^{d}$ (56)
subject to the non-slip condition that $\varPsi(x,t)=0$ for $x=(x_{1},\cdots,x_{d})$ with $x_{d}=0$. Then
$\varPsi(\xi,t)$ $=\int_{\mathbb{R}_{+}^{d}}\left(p_{b}(0,\eta;t,\xi)-p_{b}(0,\bar{\eta};t,\xi)\right)\varPsi_{0}(\eta)\textrm{d}\eta$
$+\int_{0}^{t}\int_{\mathbb{R}_{+}^{d}}\mathbb{P}_{b}^{\eta,0\rightarrow\xi,t}\left[1_{\{s>\gamma_{t}(X^{\eta})\}}g(\psi(s),s)\right]p_{b}(0,\eta;t,\xi)\textrm{d}\eta\textrm{d}s$ (57)
for $t>0$ and $\xi\in\mathbb{R}_{+}^{d}$, where $\varPsi_{0}=\varPsi(\cdot,0)$, $\zeta(\psi)=\inf\{s:\psi(s)\notin\mathbb{R}_{+}^{d}\}$ and $\gamma_{t}(\psi)$ denotes $\sup\{s\in(0,t):\psi(s)\in\mathbb{R}_{+}^{d}\}$ respectively.
Proof.
Let $D=\mathbb{R}_{+}^{d}$. The proof relies on the duality Lemma 8, established in [14]. Let $T>0$ and $\xi\in D$ be any but fixed. Let $b^{T}(x,t)=b(x,(T-t)^{+})$, and $Y$ denote the weak solution to SDE
$\textrm{d}Y_{t}=-b(Y_{t},T-t)\textrm{d}t+\sqrt{2\nu}\textrm{d}B_{t},\quad Y_{0}=\xi$ (58)
whose infinitesimal generator is $L_{-b^{T}}$. Then
$Y_{t}=\xi-\int_{0}^{t\wedge T}b(Y_{s},T-s)\textrm{d}s+\sqrt{2\nu}\int_{0}^{t\wedge T}\textrm{d}B_{s}\quad\textrm{ for all }t\geq 0.$ (59)
Let $T_{\xi}=\inf\left\{t\geq 0:Y_{t}\notin D\right\}$. Then
$Y_{t\wedge T_{\xi}}=\xi-\int_{0}^{t}1_{\left\{s<T\wedge T_{\xi}\right\}}b(Y_{s},T-s)\textrm{d}s+\sqrt{2\nu}\int_{0}^{t}1_{\left\{s<T\wedge T_{\xi}\right\}}\textrm{d}B_{s}$ (60)
for all $t\geq 0$. Let
$Z_{t}=\varPsi(Y_{t\wedge T_{\xi}},T-t)=1_{\{t<T_{\xi}\}}\varPsi(Y_{t},T-t)\quad\textrm{ for }t\leq T,$
where the second equality follows from the non-slip condition: $\varPsi$ vanishes along the boundary $\partial D$. According to ItÃŽ’s formula,
$\textrm{d}Z_{t}=\nabla\varPsi(Y_{t\wedge T_{\xi}},T-t)\cdot\textrm{d}Y_{t\wedge T_{\xi}}-\frac{\partial\varPsi}{\partial t}(Y_{t\wedge T_{\xi}},T-t)\textrm{d}t+\nu 1_{\left\{t<T\wedge T_{\xi}\right\}}\Delta\varPsi(Y_{t},T-t)\textrm{d}t.$
Using the non-slip condition again: $\varPsi$ vanishes identically on $\partial D$, so that
$\frac{\partial\varPsi}{\partial t}(Y_{t\wedge T_{\xi}},T-t)=1_{\left\{t<T_{\xi}\right\}}\frac{\partial\varPsi}{\partial t}(Y_{t},T-t)\quad\textrm{ for }t\leq T.$
Therefore
$\textrm{d}Z_{t}$ $=\sqrt{2\nu}1_{\left\{t<T_{\xi}\right\}}\nabla\varPsi(Y_{t},T-t)\cdot\textrm{d}B_{t}$
$+1_{\left\{t<T_{\xi}\right\}}\left(\nu\Delta-b\cdot\nabla-\frac{\partial}{\partial t}\right)\varPsi(Y_{t},T-t)\textrm{d}t$ (61)
for $t\leq T$. Since $\varPsi$ solves the parabolic equations, so that
$Z_{T\wedge T_{\xi}}$ $=\varPsi(\xi,T)-\int_{0}^{T\wedge T_{\xi}}1_{\{t<T_{\xi}\}}g(Y_{t},T-t)\textrm{d}t$
$+\sqrt{2\nu}\int_{0}^{T\wedge T_{\xi}}1_{\{t<T_{\xi}\}}\nabla\varPsi(Y_{t},T-t)\cdot\textrm{d}B.$ (62)
Taking expectation both sides we obtain
$\varPsi(\xi,T)=\mathbb{E}\left[\varPsi(Y_{T},0)1_{\{T<T_{\xi}\}}\right]+\mathbb{E}\left[\int_{0}^{T}1_{\{t<T_{\xi}\}}g(Y_{t},T-t)\textrm{d}t\right].$ (63)
The first term $J_{1}=\mathbb{E}\left[\varPsi_{0}(Y_{T})1_{\{T<T_{\xi}\}}\right]$ on the right hand side may be written as
$J_{1}$ $=\int_{D}\varPsi_{0}(\eta)p_{-b^{T}}^{D}(0,\xi;T,\eta)\textrm{d}\eta$
$=\int_{D}\varPsi_{0}(\eta)\left(p_{-b^{T}}(0,\xi;T,\eta)-p_{-b^{T}}(0,\xi;T,\bar{\eta})\right)\textrm{d}\eta$
$=\int_{D}\left(p_{b}(0,\eta;T,\xi)-p_{b}(0,\bar{\eta};T,\xi)\right)\varPsi_{0}(\eta)\textrm{d}\eta.$
The second term $J_{2}=\mathbb{E}\left[\int_{0}^{T}g(Y_{t},T-t)1_{\{t<T_{\xi}\}}\textrm{d}t\right]$ on the right hand side can be treated similarly
$J_{2}$ $=\int_{0}^{T}\int_{\mathbb{R}^{3}}\mathbb{E}\left[\left.1_{\{t<T_{\xi}\}}g(Y_{t},T-t)\right|Y_{T}=\eta\right]p_{-b_{T}}(0,\xi;T,\eta)\textrm{d}t$
$=\int_{0}^{T}\int_{\mathbb{R}^{3}}\mathbb{P}_{-b^{T}}^{\xi,0\rightarrow\eta,T}\left[1_{\{t<\zeta(\psi)\}}g(\psi(t),T-t)\right]p_{b}(0,\eta;T,\xi)\textrm{d}\eta\textrm{d}t$
$=\int_{0}^{T}\int_{\mathbb{R}^{3}}\mathbb{P}_{b}^{\eta,0\rightarrow\xi,T}\left[1_{\{t<\zeta(\psi\circ\tau_{T})\}}g(\psi(T-t),T-t)\right]p_{b}(0,\eta;T,\xi)\textrm{d}\eta\textrm{d}t$
$=\int_{0}^{T}\int_{\mathbb{R}^{3}}\mathbb{P}_{b}^{\eta,0\rightarrow\xi,T}\left[1_{\{T-t<\zeta(\psi\circ\tau_{T})\}}g(\psi(t),t)\right]p_{b}(0,\eta;T,\xi)\textrm{d}\eta\textrm{d}t$
where we have used the fact that, since $\nabla\cdot b=0$, $p_{-b_{T}}(0,\xi,T,\eta)$ coincides with $p_{b}(0,\eta,T,\xi)$. Therefore
$\varPsi(\xi,T)$ $=\int_{D}\left(p_{b}(0,\eta;T,\xi)-p_{b}(0,\bar{\eta};T,\xi)\right)\varPsi_{0}(\eta)\textrm{d}\eta.$
$+\int_{0}^{T}\int_{\mathbb{R}^{3}}\mathbb{P}_{b}^{\eta,0\rightarrow\xi,T}\left[1_{\{T-t<\zeta(\psi\circ\tau_{T})\}}g(\psi(t),t)\right]p_{b}(0,\eta;T,\xi)\textrm{d}\eta\textrm{d}t$ (64)
By using a similar but much simpler proof, we also have the following representation.
Theorem 10.
Let $\varPsi$ be a solution to the parabolic equation
$\left(L_{-u}-\frac{\partial}{\partial t}\right)\varPsi+g=0\quad\textrm{ in }\mathbb{R}^{d}.$ (65)
Then
$\varPsi(\xi,t)$ $=\int_{\mathbb{R}^{d}}p_{b}(0,\eta;t,\xi)\varPsi(\eta,0)\mathrm{d}\eta$
$+\int_{0}^{t}\int_{\mathbb{R}^{d}}\mathbb{P}_{b}^{\eta,0\rightarrow\xi,t}\left[f(X_{s},s)\right]p_{b}(0,\eta;t,\xi)\mathrm{d}\eta\mathrm{d}s$ (66)
for every $\xi\in\mathbb{R}^{d}$.
Data Availability Statement
The data that support the findings of this study are available from the corresponding author upon reasonable request.
Declaration of Interests
The authors report no conflict of interest.